OpenStudy (anonymous):

2(ln(x)) + ln(x-1)= ln(2) solve

OpenStudy (anonymous):

\[2\left( \ln(x) + \ln(x-1) =\ln(2) \right)\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well the 2 is outside of the brackets.. soo.. \(2ln(x) + 2ln(x-1) = ln(2)\)

OpenStudy (anonymous):

ok so then do I combine them into a multiplication problem?

OpenStudy (anonymous):

wait you can use log laws; no need to expand it

OpenStudy (anonymous):

can you show me please?

OpenStudy (anonymous):

would the 2ln just cancel?

OpenStudy (anonymous):

\[2\left(\ln(x)(x-1\right) =\ln(2)\]

OpenStudy (anonymous):

what is the difficule!! :it amounts to solve x^2-x-sqrt(2)=0

OpenStudy (anonymous):

hmm \[\ln(x)(x-1)=\frac{\ln(2)}{2} \]

OpenStudy (anonymous):

but its lohs where didyou get the x^2 from?

OpenStudy (anonymous):

logs*

OpenStudy (anonymous):

is the answer 1+ root5/2 and 1-root5/2?

OpenStudy (anonymous):

we have ln[(x-1)x]=ln(sqrt(2)) we apply the funtion e^x on both so you wil have x^2-x-sqrt(2)=0

OpenStudy (anonymous):

how did you get the sqrt(2)

OpenStudy (anonymous):

o no you were right i was just wondering how you get the part with the ln(sqrt2)?

OpenStudy (anonymous):

it's a theorem ln(x^a)=a*ln(x)

OpenStudy (anonymous):

do you understand

OpenStudy (anonymous):

yes so in this case it you took the 2 (ln(x)) and made it 2*ln(x)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

why did it end up on the other side?

OpenStudy (anonymous):

i don't understand ,you mean why i didn't finish the calcul !!!!